Solve the system of simultaneous equations, using the method of determinants (Cramer’s Rule):
{2x+y+z=0
{4x+3y+2z=2
{2x-y-3z=0

Answer by ikleyn(52778)

.
Solve the system of simultaneous equations, using the method of determinants (Cramer’s Rule):
2x +  y + z = 0
4x + 3y+ 2z = 2
2x -  y -3z = 0
~~~~~~~~~~~~~~~~~

        Solution

Your matrix

        X1	X2	X3	b
1	2	1	1	0
2	4	3	2	2
3	2	-1	-3	0


Write down the main matrix and find its determinant

        X1	X2	X3
1	2	1	1
2	4	3	2
3	2	-1	-3

Δ = -8


Replace the 1st column of the main matrix with the solution vector and find its determinant

        X1	X2	X3
1	0	1	1
2	2	3	2
3	0	-1	-3

Δ1 = 4


Replace the 2nd column of the main matrix with the solution vector and find its determinant

        X1	X2	X3
1	2	0	1
2	4	2	2
3	2	0	-3

Δ2 = -16


Replace the 3rd column of the main matrix with the solution vector and find its determinant

        X1	X2	X3
1	2	1	0
2	4	3	2
3	2	-1	0

Δ3 = 8


x = Δ1 / Δ = 4 / (-8) = -1/2
y = Δ2 / Δ = (-16) / (-8) = 2
z = Δ3 / Δ = 8 / (-8) = -1


Solution set:

x = -1/2
y = 2
z = -1

Solved.